OpenStudy (anonymous):

Will give medal Part 1. Using the two functions listed below, insert numbers in place of the letters a, b, c, and d so that f(x) and g(x) are inverses. f(x)= x+a b g(x)=cx−d Part 2. Show your work to prove that the inverse of f(x) is g(x). Part 3. Show your work to evaluate g(f(x)). Part 4. Graph your two functions on a coordinate plane. Include a table of values for each function. Include 5 values for each function. Graph the line y = x on the same graph.

OpenStudy (anonymous):

I already have the first part

OpenStudy (anonymous):

f(x) = \[\frac{ x+2 }{ 6}\]

OpenStudy (anonymous):

g(x)=6x-2

OpenStudy (adi3):

is this algebra

OpenStudy (anonymous):

Yes

OpenStudy (perl):

so you want the inverse of f(x) >?

OpenStudy (perl):

ok so you want help with part 2)

OpenStudy (anonymous):

Yeah

OpenStudy (perl):

to show that g(x) is the inverse of f(x) you have to show that f(g(x)) = x

OpenStudy (anonymous):

all of those parentheses confuse me lol.

OpenStudy (perl):

$$\Large { f(x) = (x + 2) / 6 \\ g(x) = 6x-2\\ f(g(x)) = f(6x-2) = \frac{(6x-2)+2}{6} \\ =\frac{6x}{6} = \frac{x}{1}= x }$$

OpenStudy (anonymous):

so did you set x equal to 6x-2 or something?

OpenStudy (anonymous):

wait nevermind that wouldnt make any sense. how did you get the f(g(x)) equation?

OpenStudy (perl):

What i was doing was, im testing to see if f(g(x)) = x

OpenStudy (anonymous):

Ohhh okay. So is that how you prove that the inverse of f(x) is equal to the inverse of g(x)? or is that how you evaluate f(g(x))

OpenStudy (perl):